OpenStudy (anonymous):

Simplify: (h2)5 (1 point) (0 pts) h7 I thought it was this one (0 pts) h5 (0 pts) h2 (1 pt) h10 It was this one obviously c: Can someone explain how I should of gotten h10?

OpenStudy (anonymous):

\[(h^2)^5 \] Is this it?

OpenStudy (anonymous):

The problem yesh c:

OpenStudy (anonymous):

Well let's say we have \[(x)^3\] we can rewrite it as \[(x)*(x)*(x)\] right?

OpenStudy (anonymous):

Apply that same thing to this problem.

OpenStudy (anonymous):

\[(h^2)^5= (h^2)*(h^2)*(h^2)*(h^2)*(h^2) \]

OpenStudy (anonymous):

Which can further be simplified to\[(h^2)*(h^2)*(h^2)*(h^2)*(h^2)= h*h*h*h*h*h*h*h*h*h=h^{10}\]

OpenStudy (anonymous):

See it?

OpenStudy (anonymous):

*Facepalm*

OpenStudy (anonymous):

I do see it now haha thanks

OpenStudy (anonymous):

All that means is that if you have the exponent of a variable that already has an exponent all you do is multiply the exponents

OpenStudy (anonymous):

In this case you had 2 and 5

OpenStudy (anonymous):

which makes 10

OpenStudy (anonymous):

I was thinking okay well if I lay out 2 h's then 5 h's I would have 7 but when you take h^2 It's not just one H like I was thinking for whatever reason XD

OpenStudy (anonymous):

bruh ur pic is lowkey scary asf